OpenStudy (_fluffeh_):

please help, will fan and medal. What is the value of the expression when y = –21?

OpenStudy (anonymous):

give me a sec

OpenStudy (_fluffeh_):

k

Nnesha (nnesha):

where is the expression ?

OpenStudy (anonymous):

its -21

OpenStudy (_fluffeh_):

http://prntscr.com/8nyzne

Nnesha (nnesha):

substitute y for -21 then simplify

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

the answer is the -21

OpenStudy (misty1212):

\[\frac{9-(-21)}{-6}\]

Nnesha (nnesha):

\[\huge\rm \frac{ 9-\color{red}{y} }{ -6} \] replace with -21 and then distribute by -1

OpenStudy (misty1212):

unlikely it is \(-21\) beware answers with no explanation, they are often wrong

OpenStudy (anonymous):

sorry its 5

OpenStudy (_fluffeh_):

im still confused on what the answer is, and just saying that its 5 doesn't help any >_<

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @misty1212 \[\frac{9-(-21)}{-6}\] \(\color{blue}{\text{End of Quote}}\) sole this to get the answer

Nnesha (nnesha):

solve*

Nnesha (nnesha):

 first multiply -1 tims -21 and then add 9